# Deep Unfolding with Approximated Computations for Rapid Optimization

### v1

@@ V-B Iterative Optimizer
${\bm{L}}^{(k+1)}$ $={\bm{L}}^{(k)}-{\eta}_{L}^{(k+1)}\bigl{(}\nabla_{{\bm{L}}}\mathcal{L}_{\rm o}({\bm{L}}^{(k)},{\bm{R}}^{(k)},{\bm{Y}}^{(k)};{\bm{X}})$
@@
${\bm{R}}^{(k+1)}$ $={\bm{R}}^{(k)}-{\eta}_{R}^{(k+1)}\bigl{(}\nabla_{{\bm{R}}}\mathcal{L}_{\rm o}({\bm{L}}^{(k)},{\bm{R}}^{(k)},{\bm{Y}}^{(k)};{\bm{X}})$
@@
where ${\eta}_{L}^{(k+1)},{\eta}_{R}^{(k+1)}$ are the scalar step-sizes hyperparameters. The objective gradients are obtained as
@@ V-C1 Unfolding RPCA
While the RPCA based on (25)-(28) can yield suitable offline processing, in many scenarios it is impractical for real-time applications. This stems from the fact that $(i)$ it requires multiple iterations to converge; and $(ii)$ each iteration involves heavy computations, particularly when the rank of $\bm{V}$} is not too small. Specifically, the complexity per iteration can be broken down into the following components [^1] For simplicity of presentation, we take $n\triangleq n_{1}=n_{2}$ when discussing computational complexities:
@@
Since we need to update both $\bm{L}$} and $\bm{R}$} in each iteration, the total cost of computing the low-rank is $2\mathcal{P}_{\mathrm{low}}$ flops. In total, the complexity per iteration is:
@@
The work [26] proposed to facilitate operation with a limited number of iterations by treating the thresholds $\{\zeta^{(k)}\}$ and the step-sizes $\{\eta^{(k)}\}$ (which were fixed to be identical for both $\bm{L}$} and $\bm{R}$}) as learnable parameters of an unfolded networks. This approach enables efficient operation with a fixed number of iterations. However, the per-iteration complexity remains unchanged and is still given by (31). Notably, the successful application of deep unfolding to these iterative updates motivates further reducing complexity by integrating deliberate approximation based on our methodology.
@@ V-C2 Learned Approximated RPCA
    Element-wise step-sizes - To provide additional degrees of freedom for coping with errors induced by approximations, we use matrix-valued steps-sizes, such that $\Theta_{k}=[\bm{\eta}_{L}^{(k)},\bm{\eta}_{R}^{(k)}]$, with $\bm{\eta}_{L}^{(k)}\in\mathbb{R}^{n_{1}\times r}$ and $\bm{\eta}_{R}^{(k)}\in\mathbb{R}^{n_{2}\times r}$.
@@
Init : Step sizes $\{\bm{\eta}_{L}^{(k)},\bm{\eta}_{R}^{(k)}\}$; thresholds $\{\zeta^{(k)}\}$;
@@ 2 Compute ${\bm{U}},{\bm{\Sigma}},{\bm{V}}]\leftarrow\mathrm{SVD}_{r}({\bm{X}}-{\bm{Y}}^{(0)})$;
6 if $k\in\mathcal{K}_{L}^{\rm approx}$ then
@@ 11      Set ${\bm{L}}^{(k+1)}\leftarrow{\bm{L}}^{(k)}$;
13 if $k\in\mathcal{K}_{R}^{\rm approx}$ then
@@ 18      Set ${\bm{R}}^{(k+1)}\leftarrow{\bm{R}}^{(k)}$;
LARPCA realizes an iterative optimizer with few iterations and approximated low-complexity operations, while leveraging data to mitigate performance reduction. In particular, the numbers of gradients (30), (29), and the normalized factors ($({\bm{R}}_{k}^{\top}{\bm{R}}_{k})^{-1}$ and $({\bm{L}}_{k}^{\top}{\bm{L}}_{k})^{-1}$) which require computing decreased by the following formula:
@@
We note that when no approximation is applied, i.e., $|\mathcal{K}_{R}^{\rm approx}|=|\mathcal{K}_{L}^{\rm approx}|=0$, then the per-iteration complexity in (32) coincides with that in (31). The reduction in (32) is shown in Subsection V-D to have a significant impact on latency, and the increased abstractness also facilitates coping with the mismatches induced by such deliberate approximations.
@@ V-C3 Training
For unsupervised data, i.e., when there is no ground-truth low-rank component and one only has access to a data of the form $\mathcal{D}_{\mathrm{train}}=\{{\bm{X}}^{(i)}\}_{i=1}^{|\mathcal{D}|}$, we formulate a training loss that balances relative reconstruction error and sparsity on $\hat{{\bm{Y}}}$. Specifically, since the objective in (24) does not include an implicit sparse term, we employ the unsupervised loss of (10) with the following surrogate objective
@@ V-D Experimental Study
In this section, we evaluate LARPCA in an experimental study, aiming to demonstrate our ability to achieve comparable results to non-approximated solvers, while reducing computational cost and runtime [^2] The source code and trained models for both cases studies are available at https://github.com/dviravra/Unfolded_Approximated_Optimization. We thus use the following benchmarks
@@
To assess both accuracy and efficiency trade-offs, we use two datasets: (i) Synthetic data simulated from the model in (23), where the matrices $\bm{V}$$^{\star}$ and $\bm{Y}$$^{\star}$ are generated from a Gaussian distribution complying with the low-rank and sparse priors. This synthetic setting lets us isolate the pure reduction in floating-point operations in a controlled environment that fully complies with the modeling assumptions of RPCA; and (ii) Video decomposition, using the VIRAT video dataset [50], where the low-rank component represents the static aspects of the video, while the sparse component models moving objects. This data allows us to evaluate our methodology in a setting whose resolution and temporal redundancy provide a realistic workload for measuring end-to-end runtime.
@@ V-D1 Synthetic Data
To evaluate the reduction in computations, we run each of the optimizers to reach a target error of $10^{-7}$. As summarized in Table III, LARPCA reached this error level in only $16$ iterations, while performing merely $16$ gradient updates in total. The non-approximated unfolded LRPCA required $24$ iterations (with $48$ gradient updates). This corresponds to skipping 32 out of 48 gradient steps (around $67\%$ reduction) in this scenario. Consequently, the LARPCA solver only needed about $1.76\times 10^{8}$ floating-point operations (flops) to reach the solution, which is less than half of the $3.86\times 10^{8}$ flops required by LRPCA. To reach the same error of $\approx 10^{-7}$, using standard (non-learned) optimization with fixed hyperparameters requires $6000$ iterations and $12000$ gradient calculations, resulting in around $1\times 10^{11}$ flops, which is over three orders of magnitude compared to LARPCA.
@@
As a further evaluation of our proposed method, we examined its robustness to higher-rank scenarios, which pose a significant challenge to classical LRPCA methods. Fig. 5 illustrates a comparison between the convergence profiles of LRPCA and LARPCA for data matrices of increased rank ($r=50$). We also included a baseline performance of LRPCA with a lower rank ($r=5$) as a reference. It is clearly observed in Fig. 5 that LRPCA significantly struggles when the rank of the underlying low-rank component increases. Specifically, when increasing the rank from $r=5$ to $r=50$, LRPCA does not have enough abstractness in its parameterization to efficiently minimize the reconstruction error, plateauing at a significantly higher error level. In contrast, our proposed LARPCA, despite deliberately incorporating computational approximations (in particular, approximating updates at iterations 6-9), achieves a much lower error, rapidly converging towards the desired solution. This demonstrates that the enhanced flexibility provided by learnable approximations and individual step-size matrices not only reduces computational complexity but also enables RPCA with challenging, higher-rank data.
@@
We further evaluated the effect of varying the approximation percentage on the reconstruction error. As shown in Table IV, multiple simulations were conducted at different approximation levels. The results indicate that even substantial reductions in computational complexity through approximation have only a minor impact on the relative reconstruction error, demonstrating the robustness of LARPCA to deliberate approximations.
% Approximated Error
6.25% $9.07\times 10^{-8}$
12.50% $7.75\times 10^{-8}$
25.00% $7.75\times 10^{-8}$
31.25% $8.97\times 10^{-8}$
40% $1.31\times 10^{-7}$
50.00% $1.55\times 10^{-7}$
@@ V-D2 Video Decomposition
To further validate our learned approximated optimization approach, we evaluated its effectiveness using real-world video data from the VIRAT dataset. This dataset is widely adopted for video surveillance analysis, featuring complex scenes with sparse moving objects over largely static backgrounds, which can be viewed as the low-rank and sparse structure of the RPCA model. For each experiment, we used $16$ videos for training, and $2$ for test.
Following the pre-processing approach of [26], we initially converted the color frames to grayscale, significantly reducing computational overhead while retaining essential structural information. Subsequently, we uniformly subsampled the frame rate by a factor of two to further reduce data dimensionality and computational complexity, while preserving temporal coherence. The experimental study comprised sequences from the VIRAT dataset resized to three distinct resolutions, i.e., $640\times 360$, $480\times 270$, and $320\times 180$ pixels. Each video was vectorized column-wise to form a data matrix suitable for RPCA decomposition.
The VIRAT dataset is unlabeled for RPCA decomposition, therefore, we employed unsupervised learning via (33) for training the unfolded optimizers. Due to the computationally intensive nature of this data, particularly at high resolutions, we limited the learned optimizers to operate with $K=5$ iterations, while assuming $r=2$. The reported error is evaluated as the objective
$\text{Error}(\hat{\mathbf{V}},\mathbf{X})=\frac{\|\mathbf{X}-\hat{\mathbf{V}}\|_{F}^{2}}{n_{1}\cdot n_{2}\cdot\|\mathbf{X}\|_{F}}.$ (34)
All runtime values are reported as wall-clock times on the same platform, a Lenovo IdeaPad Slim 5 14IRL8 with an Intel 13th-Gen Core i5-13420H (8C/12T, 2.10 GHz base), 16 GB RAM. In this case study, we only evaluate the learned optimizers, as the classical method with fixed hyperparameters was unstable and practically infeasible in terms of convergence time and computational cost for this video data.
Our proposed method (LARPCA) achieves approximately $40\%-58\%$ runtime reduction while preserving the same final accuracy. Table V reports quantitative results for representative sequences, clearly indicating that LARPCA achieves similar reconstruction accuracy (measured via relative squared-error per pixel) compared to the original LRPCA, while significantly reducing computational time. We also include four representative results in Fig. 6, which demonstrates the effectiveness of LARPCA in maintaining accurate foreground-background separation, despite employing approximated computational strategies.
@@
TABLE V: Comparison of Learned RPCA with and without approximations for video decomposition.
@@ VI Conclusions
We introduced a novel framework for learned approximated optimization, which extends deep unfolding by integrating approximated computations into iterative solvers to simultaneously reduce the number and complexity of iterations. By deliberately replacing computationally intensive operations with lightweight surrogates and learning to compensate for their induced mismatches, our approach enables rapid and efficient optimization suited for latency-sensitive applications. We instantiated the proposed methodology in two representative case studies of hybrid beamforming and RPCA. Our case studies consistently demonstrated the effectiveness of jointly optimizing iteration depth and per-iteration operations, paving the way for practical deployment of principled optimization methods in real-time systems.
@@ A-A Proof of Proposition 2
This proof follows the structure of conventional gradient descent convergence arguments [32, Theorem 3.6]. Define ${\bm{g}}^{(k)}\triangleq\nabla\mathcal{L}_{\rm o}({\bm{s}}^{(k)};{\bm{x}})$ and the gradient approximation error as
@@
$\bm{e}^{(k)}\triangleq\begin{cases}\tilde{{\bm{g}}}^{(k)}({\bm{s}}^{(k)};{\bm{x}})-{\bm{g}}^{(k)},&\text{if }k\in\mathcal{K}^{\rm approx},\\ \bm{0},&\text{otherwise}.\end{cases}$ (A-A.1)
Then, the update rule in (7) can be written as
${\bm{s}}^{(k+1)}={\bm{s}}^{(k)}-\bm{\eta}_{k}\odot\left(\nabla\mathcal{L}_{\rm o}({\bm{s}}^{(k)};{\bm{x}})+\bm{e}^{(k)}\right).$ (A-A.2)
We expand the recursion as follows:
$\|{\bm{s}}^{(k+1)}-{\bm{s}}^{\star}\|^{2}=\left\|{\bm{s}}^{(k)}-{\bm{s}}^{\star}-\bm{\eta}_{k}\odot\left(\bm{g}^{(k)}+\bm{e}^{(k)}\right)\right\|^{2}$
$=\|{\bm{s}}^{(k)}-{\bm{s}}^{\star}\|^{2}-2\left\langle\bm{\eta}_{k}\odot\bm{g}^{(k)},{\bm{s}}^{(k)}-{\bm{s}}^{\star}\right\rangle+\|\bm{\eta}_{k}\odot\bm{g}^{(k)}\|^{2}$
$\quad-2\left\langle\bm{\eta}_{k}\odot\bm{e}^{(k)},{\bm{s}}^{(k)}-{\bm{s}}^{\star}\right\rangle+\|\bm{\eta}_{k}\odot\bm{e}^{(k)}\|^{2}$
$\quad+2\left\langle\bm{\eta}_{k}\odot\bm{g}^{(k)},\bm{\eta}_{k}\odot\bm{e}^{(k)}\right\rangle.$ (A-A.3)
Now, by the strong convexity of $\mathcal{L}_{\rm o}$ (see  [32, Lemma 2.14]), we have
@@
$\left\langle\bm{g}^{(k)},{\bm{s}}^{(k)}-{\bm{s}}^{\star}\right\rangle\geq\mu\|{\bm{s}}^{(k)}-{\bm{s}}^{\star}\|^{2}.$ (A-A.4)
In addition, by the L-smoothness [32, Definition 2.24] of $\mathcal{L}{\rm o}$ and the fact that $\nabla\mathcal{L}{\rm o}({\bm{s}}^{\star})=0$, we have
$\|\bm{g}^{(k)}\|\leq L\|{\bm{s}}^{(k)}-{\bm{s}}^{\star}\|.$ (A-A.5)
Applying Cauchy–Schwarz inequality and the approximation error bound from (8), we get
$\left|\left\langle\bm{\eta}_{k}\odot\bm{e}^{(k)},{\bm{s}}^{(k)}-{\bm{s}}^{\star}\right\rangle\right|\leq\|\bm{\eta}_{k}\odot\bm{e}^{(k)}\|\cdot\|{\bm{s}}^{(k)}-{\bm{s}}^{\star}\|$
$\leq\delta_{k}\|{\bm{s}}^{(k)}-{\bm{s}}^{\star}\|.$ (A-A.6)
Similarly, using Cauchy–Schwarz inequality, the L-smoothness from (A-A.5), and the approximation error bound from (8), we get
$\left\langle\bm{\eta}_{k}\odot\bm{g}^{(k)},\bm{\eta}_{k}\odot\bm{e}^{(k)}\right\rangle\leq\|\bm{\eta}_{k}\odot\bm{g}^{(k)}\|\cdot\|\bm{\eta}_{k}\odot\bm{e}^{(k)}\|$
$\leq\gamma L\|{\bm{s}}^{(k)}-{\bm{s}}^{\star}\|\cdot\delta_{k},$ (A-A.7)
where we used the element-wise upper bound on $\bm{\eta}_{k}$, $\gamma$ defined before (8).
We now substitute all these bounds into the recursion in (A-A.3). To handle the mixed terms involving $\|{\bm{s}}^{(k)}-{\bm{s}}^{\star}\|\cdot\delta_{k}$, we apply Young’s inequality, which implies
$\delta_{k}\|{\bm{s}}^{(k)}-{\bm{s}}^{\star}\|\leq\frac{\gamma\mu}{2}\|{\bm{s}}^{(k)}-{\bm{s}}^{\star}\|^{2}+\frac{\delta_{k}^{2}}{2\gamma\mu}$ (A-A.8)
and
$\gamma L\delta_{k}\|{\bm{s}}^{(k)}-{\bm{s}}^{\star}\|\leq\frac{\gamma\mu}{2}\|{\bm{s}}^{(k)}-{\bm{s}}^{\star}\|^{2}+\frac{(\gamma L)^{2}}{2\gamma\mu}\delta_{k}^{2}.$ (A-A.9)
Substituting (A-A.4), (A-A.5), (A-A.8), and (A-A.9) into (A-A.3), we obtain
$\|{\bm{s}}^{(k+1)}-{\bm{s}}^{\star}\|^{2}\leq(1-2\gamma\mu+\gamma^{2}L^{2})\|{\bm{s}}^{(k)}-{\bm{s}}^{\star}\|^{2}+c_{k}\delta_{k}^{2},$ (A-A.10)
where $c_{k}$ is a constant depending on $\gamma$, $L$, and $\mu$ that collects the contribution of the approximation error terms. In particular, for $k\notin\mathcal{K}^{\rm approx}$, we have $\bm{e}^{(k)}=\bm{0}$ by definition. As a result, all error-related terms in (A-A.3) vanish, and we have
$c_{k}=0,\quad\text{for all }k\notin\mathcal{K}^{\rm approx}.$
Choosing $\gamma<1/L$ ensures that $1-\gamma\mu>0$, and unrolling the recursion gives the result in (2). It should be noted that the exact values of $c_{k}$ for $k\in\mathcal{K}^{\rm approx}$ are not required, as they can be absorbed into the definition of $\delta_{k}^{2}$, ensuring that the final bound retains the form in (2).

### v2

@@ V-B Iterative Optimizer
${\bm{L}}^{(k+1)}$ $={\bm{L}}^{(k)}-{\bm{\eta}}_{L}^{(k+1)}\odot\bigl(\nabla_{{\bm{L}}}\mathcal{L}_{\rm o}({\bm{L}}^{(k)},{\bm{R}}^{(k)},{\bm{Y}}^{(k)};{\bm{X}})$
@@
${\bm{R}}^{(k+1)}$ $={\bm{R}}^{(k)}-{\bm{\eta}}_{R}^{(k+1)}\odot\bigl(\nabla_{{\bm{R}}}\mathcal{L}_{\rm o}({\bm{L}}^{(k)},{\bm{R}}^{(k)},{\bm{Y}}^{(k)};{\bm{X}})$
@@
where ${\bm{\eta}}_{L}^{(k+1)}={\eta}_{L}^{(k+1)}{\bm{1}},{\bm{\eta}}_{R}^{(k+1)}={\eta}_{R}^{(k+1)}{\bm{1}}$ are the scalar step-sizes hyperparameters. The objective gradients are obtained as
@@ V-C1 Unfolding RPCA
While the RPCA based on (26)-(29) can yield suitable offline processing, in many scenarios it is impractical for real-time applications. This stems from the fact that $(i)$ it requires multiple iterations to converge; and $(ii)$ each iteration involves heavy computations, particularly when the rank of $\bm{V}$} is not too small. Specifically, the complexity per iteration can be broken down into the following components [^1] For brevity, we take $n\triangleq n_{1}=n_{2}$ when discussing complexities:
@@
Since we need to update both $\bm{L}$} and $\bm{R}$} in each iteration, the total cost of computing the low‑rank is $2\mathcal{P}_{\mathrm{low}}$ flops. In total, the complexity per iteration is:
@@
The work [27] proposed to facilitate operation with a limited number of iterations by treating the thresholds $\{\zeta^{(k)}\}$ and the step-sizes $\{\eta^{(k)}\}$ (which were fixed to be identical for both $\bm{L}$} and $\bm{R}$}) as learnable parameters of an unfolded network. This approach enables efficient operation with a fixed number of iterations. However, the per-iteration complexity remains unchanged and is still given by (32). Notably, the successful application of deep unfolding to these iterative updates motivates further reducing complexity by integrating deliberate approximation based on our methodology.
@@ V-C2 Learned Approximated RPCA
    Element-wise step-sizes - To provide additional degrees of freedom for coping with errors induced by approximations, we use matrix-valued step-sizes, such that $\Theta_{k}=[\bm{\eta}_{L}^{(k)},\bm{\eta}_{R}^{(k)}]$, with $\bm{\eta}_{L}^{(k)}\in\mathbb{R}^{n_{1}\times r}$ and $\bm{\eta}_{R}^{(k)}\in\mathbb{R}^{n_{2}\times r}$.
@@
Init : Step-sizes $\{\bm{\eta}_{L}^{(k)},\bm{\eta}_{R}^{(k)}\}$; thresholds $\{\zeta^{(k)}\}$;
@@ 2 Compute ${\bm{U}},{\bm{\Sigma}},{\bm{V}}]\leftarrow\mathrm{SVD}_{r}({\bm{X}}-{\bm{Y}}^{(0)})$;
6 if $k\notin\mathcal{K}_{L}^{\rm approx}$ then
@@ 11      Set ${\bm{L}}^{(k+1)}\leftarrow{\bm{L}}^{(k)}$;
13 if $k\notin\mathcal{K}_{R}^{\rm approx}$ then
@@ 18      Set ${\bm{R}}^{(k+1)}\leftarrow{\bm{R}}^{(k)}$;
LARPCA realizes an iterative optimizer with few iterations and approximated low-complexity operations. In particular, the numbers of computed gradients (31), (30), and the normalized factors ($({\bm{R}}_{k}^{\top}{\bm{R}}_{k})^{-1}$ and $({\bm{L}}_{k}^{\top}{\bm{L}}_{k})^{-1}$), decreased by the following formula:
@@
When no approximation is applied, i.e., $|\mathcal{K}_{R}^{\rm approx}|=|\mathcal{K}_{L}^{\rm approx}|=0$, then the per-iteration complexity in (33) coincides with (32). The reduction in (33) is shown in Subsection V-D significantly enhance latency, and the increased abstractness facilitates coping with the mismatches induced by such deliberate approximations.
@@ V-C3 Training
For unsupervised data, i.e., when there is no ground-truth low-rank component and one only has access to $\mathcal{D}_{\mathrm{train}}=\{{\bm{X}}^{(i)}\}_{i=1}^{|\mathcal{D}|}$, we formulate a training loss that balances relative reconstruction error and sparsity on $\hat{{\bm{Y}}}$. Specifically, since the objective in (25) does not include an implicit sparse term, we employ the unsupervised loss of (10) with the following surrogate objective
@@ V-D Experimental Study
In this section, we evaluate LARPCA in an experimental study, aiming to demonstrate our ability to achieve comparable results to non-approximated solvers, while reducing computational cost and runtime [^2] The source code and trained models for both case studies are available at https://github.com/dviravra/Unfolded_Approximated_Optimization. We thus use the following benchmarks
@@
To assess accuracy and efficiency, we consider: (i) Synthetic data simulated from the model in (24), where the matrices $\bm{V}$$^{\star}$ and $\bm{Y}$$^{\star}$ are generated from a Gaussian distribution complying with the low-rank and sparse priors. This synthetic setting lets us isolate the reduction in flops in a controlled environment that fully complies with the modeling assumptions of RPCA; and (ii) Video decomposition, using the VIRAT video dataset [46], where the low-rank component represents the static aspects of the video, while the sparse component models moving objects.
@@ V-D1 Synthetic Data
To evaluate the reduction in computations, we run each of the optimizers to reach a target error of $10^{-7}$. As summarized in Table III, LARPCA reached this error level in only $16$ iterations, while performing merely $16$ gradient updates in total. The non-approximated unfolded LRPCA required $24$ iterations (with $48$ gradient updates). This corresponds to skipping 32 out of 48 gradient steps (around $67\%$ reduction). Consequently, LARPCA only needed $1.76\times 10^{8}$ flops to reach the solution, which is less than half of the $3.86\times 10^{8}$ flops required by LRPCA. To reach the same error of $\approx 10^{-7}$, using standard (non-learned) optimization with fixed hyperparameters requires $6000$ iterations and $12000$ gradient calculations, resulting in around $1\times 10^{11}$ flops, which is over three orders of magnitude compared to LARPCA.
@@
As a further evaluation of our proposed method, we examined its robustness to higher-rank scenarios. Fig. 5 illustrates a comparison between the convergence profiles of LRPCA and LARPCA for data matrices of increased rank ($r=50$), showing that LRPCA struggles when the rank of the underlying low-rank component increases. Specifically, when increasing the rank from $r=5$ to $r=50$, LRPCA plateaus at a significantly higher error level. LARPCA, despite incorporating approximations (at iterations 6-9), achieves a much lower error, rapidly converging towards the desired solution. This demonstrates that the enhanced flexibility provided by learnable approximations and individual step-size matrices not only reduces computational complexity but also enables RPCA with challenging, higher-rank data.
@@
We further evaluate the effect of varying the approximation percentage on the reconstruction error. As shown in Table IV, multiple simulations were conducted at different approximation levels. The results indicate that even substantial reductions in computational complexity through approximation have only a minor impact on the relative reconstruction error.
% Approximated Error % Approximated Error
6.25% $9.07\times 10^{-8}$ 31.25% $8.97\times 10^{-8}$
12.50% $7.75\times 10^{-8}$ 40% $1.31\times 10^{-7}$
25.00% $7.75\times 10^{-8}$ 50.00% $1.55\times 10^{-7}$
@@ V-D2 Video Decomposition
To further validate our learned approximated optimization approach, we evaluated its effectiveness using real-world video data from the VIRAT dataset. This dataset is widely adopted for video surveillance analysis, featuring complex scenes with sparse moving objects over largely static backgrounds, which can be viewed as the low-rank and sparse structure of the RPCA model. For each experiment, we used $16$ videos for training, and $2$ for test. Following the pre-processing approach of [27], we initially convert color frames to grayscale. Subsequently, we uniformly subsampled the frame rate by a factor of two to further reduce data dimensionality and computational complexity, while preserving temporal coherence. The experimental study comprised sequences from the VIRAT dataset resized to three distinct resolutions, i.e., $640\times 360$, $480\times 270$, and $320\times 180$ pixels. Each video was vectorized column-wise to form a data matrix suitable for RPCA decomposition.
@@
TABLE V: Comparison of Learned RPCA for video decomposition.
The VIRAT dataset is unlabeled for RPCA decomposition, therefore, we employed (34) for training the unfolded optimizers. Due to the computationally intensive nature of this data, particularly at high resolutions, the learned optimizers operate with $K=5$ iterations, while assuming $r=2$. The reported error is evaluated as the objective $\text{Error}(\hat{\mathbf{V}},\mathbf{X})=\frac{\|\mathbf{X}-\hat{\mathbf{V}}\|_{F}^{2}}{n_{1}\cdot n_{2}\cdot\|\mathbf{X}\|_{F}}$. All runtime values are reported as wall-clock times on the same platform, a Lenovo IdeaPad Slim 5 14IRL8 with an Intel 13th-Gen Core i5-13420H (8C/12T, 2.10 GHz base), 16 GB RAM. Here, we only evaluate the learned optimizers, as optimization with fixed hyperparameters was unstable and lengthy. LARPCA achieves approximately $40\%-58\%$ runtime reduction while preserving the same final accuracy. Table V reports quantitative results for representative sequences, clearly indicating similar reconstruction accuracy (measured via relative squared-error per pixel) compared to LRPCA, while significantly reducing computational time. We also include four representative results in Fig. 6, which demonstrates the effectiveness of LARPCA in foreground-background separation, despite employing approximated computational strategies.
@@ VI Conclusions
We introduced a novel framework for learned approximated optimization, which extends deep unfolding by integrating approximated computations into iterative solvers. By deliberately replacing computationally intensive operations with lightweight surrogates and learning to compensate for their induced mismatches, we enable rapid and efficient optimization. We instantiated the proposed methodology in two representative case studies of hybrid beamforming and RPCA, demonstrating the effectiveness of jointly optimizing iteration depth and per-iteration operations.
@@ A-A Proof of Proposition 2
For brevity, let $f({\bm{s}})\triangleq\mathcal{L}_{\rm o}({\bm{s}};{\bm{x}})$, ${\bm{g}}^{(k)}\triangleq\nabla f({\bm{s}}^{(k)})$, and ${\bm{D}}_{k}\triangleq\mathrm{diag}(\bm{\eta}_{k})$. Also define
@@
${\bm{e}}^{(k)}\triangleq\begin{cases}\tilde{{\bm{g}}}^{(k)}({\bm{s}}^{(k)};{\bm{x}})-\nabla f({\bm{s}}^{(k)}),&k\in\mathcal{K}^{\rm approx},\\ {\bm{0}},&k\notin\mathcal{K}^{\rm approx}.\end{cases}$
Then ${\bm{s}}^{(k+1)}={\bm{s}}^{(k)}-{\bm{D}}_{k}\big({\bm{g}}^{(k)}+{\bm{e}}^{(k)}\big)$ and $\|{\bm{D}}_{k}{\bm{e}}^{(k)}\|_{2}\leq\delta_{k}$.
By the $L$-smoothness of the objective, we have that
$f({\bm{s}}^{(k+1)})\leq f({\bm{s}}^{(k)})-\left\langle{\bm{g}}^{(k)},{\bm{D}}_{k}{\bm{g}}^{(k)}\right\rangle-\left\langle{\bm{g}}^{(k)},{\bm{D}}_{k}{\bm{e}}^{(k)}\right\rangle$
$\qquad+\frac{L}{2}\|{\bm{D}}_{k}{\bm{g}}^{(k)}+{\bm{D}}_{k}{\bm{e}}^{(k)}\|_{2}^{2}$
$=f({\bm{s}}^{(k)})-\left\langle{\bm{g}}^{(k)},{\bm{D}}_{k}{\bm{g}}^{(k)}\right\rangle-\left\langle{\bm{g}}^{(k)},{\bm{D}}_{k}{\bm{e}}^{(k)}\right\rangle+\frac{L}{2}\|{\bm{D}}_{k}{\bm{g}}^{(k)}\|_{2}^{2}$
$\qquad+L\left\langle{\bm{D}}_{k}{\bm{g}}^{(k)},{\bm{D}}_{k}{\bm{e}}^{(k)}\right\rangle+\frac{L}{2}\|{\bm{D}}_{k}{\bm{e}}^{(k)}\|_{2}^{2}.$ (A-A.1)
Next, since all entries of $\bm{\eta}_{k}$ lie in $\underline{\eta}_{k},\bar{\eta}_{k}]$, then
$\|{\bm{D}}_{k}{\bm{g}}^{(k)}\|_{2}^{2}$ $=({\bm{g}}^{(k)})^{\top}{\bm{D}}_{k}^{2}{\bm{g}}^{(k)}\leq\bar{\eta}_{k}({\bm{g}}^{(k)})^{\top}{\bm{D}}_{k}{\bm{g}}^{(k)}$
$=\bar{\eta}_{k}\left\langle{\bm{g}}^{(k)},{\bm{D}}_{k}{\bm{g}}^{(k)}\right\rangle.$ (A-A.2)
Substituting (A-A.2) into (A-A.1) gives
$f({\bm{s}}^{(k+1)})$ $\leq f({\bm{s}}^{(k)})-\left(1-\frac{L}{2}\bar{\eta}_{k}\right)\left\langle{\bm{g}}^{(k)},{\bm{D}}_{k}{\bm{g}}^{(k)}\right\rangle$
$-\left\langle{\bm{g}}^{(k)},{\bm{D}}_{k}{\bm{e}}^{(k)}\right\rangle$
$+L\left\langle{\bm{D}}_{k}{\bm{g}}^{(k)},{\bm{D}}_{k}{\bm{e}}^{(k)}\right\rangle+\frac{L}{2}\|{\bm{D}}_{k}{\bm{e}}^{(k)}\|_{2}^{2}.$ (A-A.3)
Applying Cauchy–Schwarz and using $\|{\bm{D}}_{k}{\bm{e}}^{(k)}\|_{2}\leq\delta_{k}$ together with $\|{\bm{D}}_{k}{\bm{g}}^{(k)}\|_{2}\leq\bar{\eta}_{k}\|{\bm{g}}^{(k)}\|_{2}$, (A-A.3) yields
$f({\bm{s}}^{(k+1)})$ $\leq f({\bm{s}}^{(k)})-\left(1-\frac{L}{2}\bar{\eta}_{k}\right)\left\langle{\bm{g}}^{(k)},{\bm{D}}_{k}{\bm{g}}^{(k)}\right\rangle$
$+(1+L\bar{\eta}_{k})\|{\bm{g}}^{(k)}\|_{2}\delta_{k}+\frac{L}{2}\delta_{k}^{2}.$ (A-A.4)
@@
Since ${\bm{D}}_{k}\succeq\underline{\eta}_{k}{\bm{I}}$, then $\left\langle{\bm{g}}^{(k)},{\bm{D}}_{k}{\bm{g}}^{(k)}\right\rangle\geq\underline{\eta}_{k}\|{\bm{g}}^{(k)}\|_{2}^{2}$ and thus
$f({\bm{s}}^{(k+1)})\leq f({\bm{s}}^{(k)})-c_{k}\|{\bm{g}}^{(k)}\|_{2}^{2}+(1+L\bar{\eta}_{k})\|{\bm{g}}^{(k)}\|_{2}\,\delta_{k}+\frac{L}{2}\delta_{k}^{2},$
where $c_{k}\triangleq\underline{\eta}_{k}(1-\frac{L}{2}\bar{\eta}_{k})$. By Young’s inequality, we have $(1+L\bar{\eta}_{k})\|{\bm{g}}^{(k)}\|_{2}\,\delta_{k}\leq\frac{c_{k}}{2}\|{\bm{g}}^{(k)}\|_{2}^{2}+\frac{(1+L\bar{\eta}_{k})^{2}}{2c_{k}}\delta_{k}^{2}$, so that
$f({\bm{s}}^{(k+1)})\leq f({\bm{s}}^{(k)})-\frac{c_{k}}{2}\|{\bm{g}}^{(k)}\|_{2}^{2}+\left(\frac{L}{2}+\frac{(1+L\bar{\eta}_{k})^{2}}{2c_{k}}\right)\delta_{k}^{2}.$
The strong convexity ($\|\nabla f({\bm{s}})\|_{2}^{2}\geq 2\mu\big(f({\bm{s}})-f({\bm{s}}^{\star})\big)$) implies
$f({\bm{s}}^{(k+1)})-f({\bm{s}}^{\star})\leq(1-\mu c_{k})\big(f({\bm{s}}^{(k)})-f({\bm{s}}^{\star})\big)+C_{k}\delta_{k}^{2},,$
where $C_{k}\triangleq\frac{L}{2}+\frac{(1+L\bar{\eta}_{k})^{2}}{2c_{k}}$. Since $c_{k}\geq c$ and $C_{k}\leq C$,
$f({\bm{s}}^{(k+1)})-f({\bm{s}}^{\star})\leq(1-\mu c)\big(f({\bm{s}}^{(k)})-f({\bm{s}}^{\star})\big)+C\delta_{k}^{2}.$ (A-A.5)
Unrolling recursion (A-A.5) over $k=0,\ldots,K-1$ yields (9).